### v2

Abstract
This paper integrates a novel concept, coined as an individual–Environmental, Social, Economics (ESE) score with a joint liability model in order to gain insights into optimal group sizes and individual–ESE score relationships. We adopt a mean–variance utility function for farmers to effectively capture risk associated with anticipated profits. Through a set of simulation exercises, the paper studies a number of implications of incorporating $ESE$ scores into credit evaluation systems, offering, in the process, a nuanced comprehension of the repercussions under various climate conditions.
Keywords: Climate Change, ESG, Joint Liability, Sustainable Agriculture, ESE, Mean-Variance Utility.
1 Introduction
In recent years the importance of incorporating sustainability factors into financial decision–making has gained a major traction. At the same time, while credit metrics for corporations have been extensively studied, there remains a noticeable gap in the assessment of individual borrowers, particularly farmers seeking credit to finance their agricultural operations. To address this gap, this paper introduces a novel concept of scoring system coined as an Environmental, Sustainability, and Economics (ESE), which considers not only traditional credit metrics, including economic factors, but importantly also environmental and social aspects which contribute to sustainable agricultural practices and which are so prominent in the standard Environmental, Sustainability, and Governance (ESG) score. See ztwp2024 for a further discussion and analyis of this individual–ESE concept.
In particular, in this paper, we propose a novel link between an individual–ESE score and a joint liability model and provide lenders and borrowers (where lenders are insurance companies and borrowers are farmers) with optimal scores and corresponding optimal group sizes by maximizing total utility of a group of individual borrowers. We start our analysis with the simplest case of two borrowers in the group. Later we generalize it to a group of $n$ individual borrowers. It is important to incorporate the individual–ESE score mechanism into the joint liability model as this can benefit both lenders and borrowers. For lenders, this mechanism provides a better understanding of the borrower’s risk profile, enabling them to make informed lending decisions and, in the process, ensure more sustainable joint liability contracts. It also provides information about the optimal group size for lending purposes. For borrowers, the mechanism provides them with a clearer picture of their ranking relative to their peers, allowing them to identify strengths and weaknesses and work on improving their individual–ESE scores. It also provides information about the number of peers needed for a group borrowing purpose, helping borrowers form groups that are more likely to succeed and have improved access to credit. Therefore we argue that incorporating the individual–ESE score mechanism into credit evaluation processes can lead to better–informed lending decisions, improved access to credit, and sustainable development outcomes.
The remaining parts of this paper are organized as follows. We introduce the set-up and assumptions of the model in Section 2. We derive a set of results from the models consisting of a group size of two and a group size of $n$ in Sections 3 and 4 respectively. In Section 5 we refine the utility functional form of the model to that of a mean variance, which incorporates both the expected value of return and risk simultaneously in the face of climate change. As the resulting model does not have clean and closed-form analytical expressions, a set of simulation exercises is provided based on a set of climate scenarios. Section 6 concludes the paper with several qualifications and avenues for future research.
2 The Framework for an Individual ESE-Joint Liability Model
We are interested in examining a microcredit lending scenario in which a benevolent lender extends loans to a group of self–selected micro–entrepreneur borrowers who are jointly liable for repayment of the loan. Although each member of the group is given an individual loan, the group is collectively responsible for repaying the loans, and if the total loan is repaid, the entire group can qualify for a subsequent loan. The loans are then invested in a set of projects, which have an equal chance of success and can be either disjoint or correlated. It is assumed that the group members know the realized output of each project, while the lender does not. Both the lender and the borrowers have to make decisions in this set up. The lender has to decide on the optimal contract structure, the degree of joint liability, and the level of punishment in the event of a strategic default, while the borrowers have to decide whether to repay or default the loan.
There are two prior studies which have inspired us in particular to write this paper. First rezaei2017optimal show that a group size can affect lending efficiency. Larger groups are more reliable in repaying loans for riskier projects, but the group size cannot become too large. Joint liability has a positive effect on borrowers’ repayment amount and welfare compared to individual liability, provided that group members impose strong punishment strategies against any strategically defaulting member. rezaei2017optimal also examine the possibility of relaxing some of the simplifying assumptions made in the basic model of joint liability. Second, in the context of climate change, castaing2021joint investigates the impact of risk–sharing networks on the adoption of weather shock management strategies in Burkina Faso; particularly, he studies how the principle of forced solidarity may reduce efforts to adopt techniques that mitigate exposure to climate change. The study finds that a system based on mutual assistance between farmers may reduce their incentives to adopt risk–mitigating strategies, both incrementally and transformationally. The study highlights the need for complementary mechanisms, such as insurance, to help farmers reduce their vulnerability to climate change.
This paper pursues two primary objectives. The first one is to determine an optimal group size that maximizes the benefit for the borrowers, while ensuring that the lender breaks even. A larger group size can have a positive effect on the expected repayment amount, as more people are liable for repaying defaulted payments, thus ensuring a higher rate of repayment. However, a larger group size can also pose a threat to repaying members, as they must repay all defaulting peers’ repayments, when everyone else in the group defaults on repayment. Therefore this paper proposes a method to determine the optimal group size that enables joint liability to achieve its maximum loan ceiling. A second objective of this paper is to derive an optimal individual–ESE score. The importance of such a scoring system for both lenders and borrowers cannot be overstated. For lending companies, having access to a specific scoring mechanism for individual borrowers can provide a better understanding of their creditworthiness and repayment prospect. This information is essential for lending companies in determining the risk associated with lending to a particular borrower. Additionally the optimal group size of borrowers can be determined using the scoring system, which enables the lending company to mitigate risk by lending to a diversified portfolio. For borrowers, having a scoring system can equip them with a clearer picture of their ranking and eligibility for loans. This information is critical for farmers as they search for a group for the borrowing purposes since this can help them determine the number of peers needed in a group to secure a loan and improve their chances of obtaining a funding. Therefore, both lenders and borrowers can benefit from the implementation of such a comprehensive scoring system.
2.1 Model Assumptions
Prior to constructing the individual ESE–joint liability model, we state a number of assumptions underlying it.
ASSUMPTION 1. The borrowers (who are the individual farmers) produce an agricultural product. For each borrower, the probability of high production (i.e., a successful project) is the same as given by $e$ and the probability of low production (i.e., a failed project) is given by $1-e$ respectively.
ASSUMPTION 2. A higher individual–ESE score means that the farmer has a stronger ability to succeed in their project due to a high-quality comprehensive evaluation. For simplicity the probability of high yield is assumed to be linearly related to the borrower’s $ESE$ score; i.e., we assume that there is a linear relationship between $e$ and $ESE$ score: $e=k\cdot ESE+b$. The parameter $k$ represents the weight or strength of the impact of the borrower’s $ESE$ score on the probability of high yield. A higher value of $k$ indicates that a higher $ESE$ score leads to a more significant positive impact on the probability of success. This aligns well with the idea that a borrower’s positive $ESE$ practices contribute to a project success. In other words, higher $ESE$ scores indicate a better alignment with responsible practices. The constant $b$ represents a baseline/constant contribution of factors beyond an individual’s control, such as climatic and geographical conditions. Even if a borrower has a lower $ESE$ score, this constant term acknowledges that external factors can still influence the project’s success. It is essential to account for such factors in order to avoid unfairly penalizing borrowers for circumstances beyond their own control.
The selection of the linear equation $e=k\cdot ESE+b$ is motivated by both the characteristics of the variables involved and its practical applicability. Notably, the non-negativity of $k$, $b$, and $ESE$ ensures that the resulting probability, denoted as $e$, maintains its intrinsic non-negative nature. As $ESE$ encompasses a continuous range from 0 to 100, it inherently aligns with this non-negative property. It is important to stress that the intuitive requirement for $e$ to lie between 0 and 1, which is a fundamental property of a probability, is elegantly met by setting $b=0$ and $k=\frac{1}{100}$. By adhering to these values, the probability of $e$ adheres to the established probability range. In addition, the model accommodates real-world scenarios by allowing for a non-zero baseline influence ($b>0$), often linked to exogenous factors. To this end, a modest value such as $b=0.3$ can be accommodated without violating the probability constraints. Notably, this adjustment merely necessitates setting $k=\frac{0.7}{100}$, thereby guaranteeing that $e$ remains within the bounds of a probability as it should be. While, at the same time, it also acknowledges that $e$ cannot span the interval from 0 to 0.3, this limitation is justified by underscoring $e$ as a representation of a probability of a high yield, which is inherently a positive entity in practical settings. Hence the rationale underlying the proposed equation effectively marries mathematical coherence with the nuanced demands from a real–world application.
ASSUMPTION 3. The amount of loan secured by each borrower is denoted as $L$, while the amount for this borrower to be repaid is denoted as $w$. The unit of a selling price of the product is denoted as $p$, and the total income of a high production is denoted as $p\overline{Y}$, while the total income of a low production is denoted as $p\underline{Y}$.
ASSUMPTION 4. High-yield borrowers repay their loans and low-yield borrowers pay back all the income they earn. As long as this group repays the total loans, this group can continue to secure subsequent loans. In the event that all borrowers are unsuccessful in their projects, the group will not be eligible for any future loans. This may result in a negative credit history for all members in the group, which could lead to further restrictions placed on them, such as being unable to use public transportation or purchase high-cost items. Therefore, if all borrowers fail, they may have to use all of their earnings to avoid receiving negative credit reports, even though their total earnings will be less than the total repayments.
ASSUMPTION 5. To increase the successful probability $e$, the borrower incurs additional costs, such as purchasing green pesticides, advanced tools, etc. A convex disutility function, denoted by $C(e)$, is included in the specification of the total utility function in the proposed model.
2.2 Joint Liability Contracts with a Group Size of Two
Under a joint liability agreement, each farmer’s ex-ante expected profit $\pi_{i}$ can be written as
$\pi_{i}=e^{2}[p\overline{Y}-w]+e(1-e)[p\overline{Y}+p\underline{Y}-2w]-C(e)$ (2.1)
We emphasize that, in this model, both farmers have a chance of achieving a successful harvest $\overline{Y}$ with a probability of $e^{2}$, which gives them a profit of $p\overline{Y}-w$. The probability of one farmer defaulting is given by $e(1-e)$, which means that farmer $i$ would receive his/her own surplus from a successful harvest minus the other farmer’s deficit, $p\overline{Y}+p\underline{Y}-2w$. If both farmers fail at the same time, they will pay all of their incomes to the lenders, which means that their profit will be zero. The detailed profit distributions are summarized conveniently in Table 1.
Case Description Probability Profit=Earning-Payment
One farmer succeeds, another farmer succeeds too. $e^{2}$ $p\overline{Y}-w$
One farmer succeeds, another farmer fails. $e(1-e)$ $p\overline{Y}+p\underline{Y}-2w$
One farmer fails, another farmer succeeds. $(1-e)e$ $p\underline{Y}-p\underline{Y}=0$
One farmer fails, another farmer fails. $(1-e)^{2}$ $p\underline{Y}-p\underline{Y}=0$
Table 1: Profit Distributions
Before setting the total objective (utility) function for the whole group in the model, we first list a number of exogenous constraints which we introduce to the corresponding optimization program.
CONSTRAINT 1. If the group wants a subsequent loan from the lender, the group’s total repayment amount must be affordable even in the worst–case scenario when everyone else’s project has failed except for him/hers:
[…]
CONSTRAINT 2. Strategic default in a joint liability contract occurs when one party (or more parties) intentionally chooses (or choose) to default on his/her obligation (or their obligations), even though he/she has (or they have) the ability to pay back the loan. This is often done in order to gain a strategic advantage, such as to force the other parties to pay a greater share of the debt or renegotiate the terms of the contract. To avoid a strategic default for each successful borrower, the payoff of strategically defaulting cannot be larger than the payoff of repaying and being refinanced.
Assuming that the other borrower also succeeds in his/her project, this borrower will only be responsible for his/her own payment. In such a scenario the earnings plus the discounted cost of the next loan would at least be equal to the earnings of strategically defaulting.
[…]
CONSTRAINT 3. If the other borrower fails in his/her project, then this borrower needs to pay for himself/herself as well as part of the other borrower’s repayment amount. Now, the earnings plus the discounted cost of the next loan would also at least equal the earnings of strategically defaulting.
[…]
CONSTRAINT 4. The lender must be able to sustain the lending game, or at least break even, over the periods. Thus the expected repayment amount of each borrower has to be at least as large as $L(1+\epsilon)$, where $\epsilon$ is a risk-free rate. It is important to note that given that all constraints have been established under the assumption that the group will successfully repay the total loans to ensure their eligibility for subsequent loan rounds, we have not factored in the scenario where all borrowers within the group experience failure on the left hand side of Constraint 4.
[…]
Since $p\underline{Y}-w<0$, it is clear that Constraint 3 is tighter than Constraint 2. In the ensuing analysis, we will only consider Constraint 1, Constraint 3, and Constraint 4. Firstly, $\pi_{i}$ and $w$ have a strictly negative relationship as can be observed from Equation (2.1). So, to maximize $\pi_{i}$, $w$ has to be set at a minimum value. From Constraint 4, the minimum value of $w$ should be set at
$w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ (2.6)
. Next we substitute $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into Constraint 1, and rewrite the equation as:
$\quad 2\cdot\frac{L(1+\varepsilon)}{1-(1-e)^{2}}\leqslant p\overline{Y}+p\underline{Y},$ (2.7)
and the first loan ceiling bound can be expressed in Equation (2.8) as
$L_{1}\leqslant\frac{p\overline{Y}+p\underline{Y}}{2(1+\varepsilon)}\left[1-(1-e)^{2}\right].$ (2.8)
After substituting $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into Constraint 3, we obtain
$p\overline{Y}\leqslant p\overline{Y}+p\underline{Y}-2\frac{L(1+\varepsilon)}{1-(1-e)^{2}}+\delta L,$ (2.9)
$\left(\frac{2(1+\varepsilon)}{1-(1-e)^{2}}-\delta\right)L\leqslant P\underline{Y},$ (2.10)
This allows us to obtain another bound for the loan ceiling as in Equation (2.11)
$L_{2}\leqslant\frac{p\underline{Y}}{\frac{2(1+\varepsilon)}{1-(1-e)^{2}}-\delta}$ (2.11)
Next we substitute $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into the objective function (which is equation (2.1) and replace $e=k\cdot E+b$. The ex-ante expected profit function of an individual borrower can now be rewritten as
$\max_{E}e^{2}(p\overline{Y}-w)+e(1-e)[p\overline{Y}+p\underline{Y}-2w]-C(e)$ (2.12)
$=e^{2}p\overline{Y}+e(1-e)(p\overline{Y}+p\underline{Y})-\left(2e-e^{2}\right)w-C(e)$
$=e^{2}p\overline{Y}+e(1-e)(p\overline{Y}+p\underline{Y})-L(1+\varepsilon)-C(e)$
$=(kE+b)^{2}\cdot p\overline{Y}+(kE+b)(1-kE-b)(p\overline{Y}+p\underline{Y})-L(1+\varepsilon)-C(e)$
To simplify the problem even further, we set $C(e)=\frac{1}{2}ce^{2}$ (where the cost parameter $c$ is used to scale the total cost) as the disutility function. The disutility function $C(e)=\frac{1}{2}ce^{2}$ is a commonly used function in economics to represent the cost or disutility associated with effort or work. See e.g. clark1996satisfaction and kahneman1997back. The function is quadratic, meaning that as an effort increases, the cost or disutility of that effort increases at an increasing rate. In the context of an individual agricultural farmer, it is reasonable to use a disutility function to acknowledge that while some effort may be necessary, there is a point at which additional effort becomes more and more unpleasant or burdensome. For instance, a farmer may be willing to put in an extra effort to cultivate crops or tend to a livestock, but as the required effort increases, the farmer may experience fatigue, stress, and reduced productivity, which can negatively impact their overall well–being. Furthermore, the quadratic form of the disutility function is also supported by empirical evidence on the relationship between effort and well–being. It is reasonable to assume that farmers will experience a phenomenon known as a diminishing marginal utility of effort, which implies that the additional unit of well-being gained by exerting more effort decreases as the level of effort increases.
2.3 Main Results
Now that we have listed the assumptions and set-up for the model, we proceed to state a number of key results expressed conveniently in the form of propositions. These propositions provide useful insights into the optimal individual–ESE scores, optimal group size, and loan ceiling constraints. PROPOSITION 1 (Optimal individual–ESE score). Finding an optimal ESE score for both the lenders and borrowers in an individual ESE–joint liability model is critical because it can help address the challenge of promoting sustainability in agricultural production while ensuring access to credit for small-scale farmers. By incorporating environmental and social factors into credit evaluation systems, it can incentivize farmers to adopt sustainable practices that can lead to positive environmental outcomes and long–term economic benefits. At the same time lenders can mitigate the risks associated with lending to farmers by incorporating sustainability factors into their credit evaluation process. To find the optimal individual–$ESE$ score (noted in formulas as $E$), we take a first-order partial derivative w.r.t $E$ of the ex-ante expected payoff function to obtain the following result
$\frac{\partial\pi}{\partial E}$ $=2(kE+b)\cdot k(p\overline{Y})+(k-2(kE+b)\cdot k)\cdot[p\overline{Y}-p\underline{Y}]-k\cdot\frac{\partial C(e)}{\partial e}$ (2.13)
$=kp\overline{Y}-[k-2(kE+b)\cdot k]p\underline{Y}-kc\cdot(kE+b)=0.$
After rearranging terms, we obtain an optimal individual-$ESE$ score (where $E$ is between 0 and 100) as
$E=\frac{p\overline{Y}+(2b-1)p\underline{Y}-cb}{ck-2kp\underline{Y}}$ (2.14)
PROPOSITION 2 (Collapsing two constraints on loan ceiling to a single constraint). The two loan ceiling constraints (Constraint 1 and Constraint 3) arise from analyzing the success or failure of the other co–loaners assuming their own success in the project, but most of the previous studies in this area have not explored the connection between these constraints. Notably, in rezaei2017optimal, they found that a feasible range of loan ceiling depends on the borrower’s discount rate $\delta$. Nevertheless, it is possible to reduce these two constraints to a single constraint in our model. Recall the two constraints on the loan ceiling in Equation (2.8) and Equation (2.11), labelled respectively as $L_{1}$ and $L_{2}$.
$L_{1}\leqslant\frac{p\overline{Y}+p\underline{Y}}{2(1+\varepsilon)}\left[1-(1-e)^{2}\right]$
$L_{2}\leqslant\frac{P\underline{Y}}{\frac{2(1+\varepsilon)}{1-(1-e)^{2}}-\delta}$
Next we show that $L_{1}$ is always larger than $L_{2}$, which means that we can only use $L_{2}$ as a final loan ceiling bound. It simplifies the problem and potentially makes it easier to analyze and solve the model. In addition, $L_{2}$ takes $\delta$ into consideration, which may include more information and additional key factors. This conclusion is illustrated in Figure 1.
Figure 1: The Relationship between Two Loan Ceiling Constraints
Proof: If $L_{1}\leqslant L_{2}$, then
$\frac{p\overline{Y}+p\underline{Y}}{2(1+\varepsilon)}\left[1-(1-e)^{2}\right]\leqslant\frac{p\underline{Y}}{\frac{2(1+\varepsilon)}{1-(1-e)^{2}}-\delta}$ (2.15)
After rearranging terms, we obtain
$\delta\geqslant\frac{\overline{Y}\cdot 2(1+\varepsilon)}{\left[1-(1-e)^{2}\right](\overline{Y}+\underline{Y})}>1$ (2.16)
From the above results, if $L_{1}\leqslant L_{2}$, the discount factor $\delta$ is always greater than 1, which leads to a contradiction. This leads us to conclude that $L_{1}>L_{2}$. PROPOSITION 3 (A higher individual–ESE score leads to a higher loan ceiling). Next we examine the relationship between the individual–ESE score and the loan ceiling. Intuitively a borrower with a high-quality ESE score should be eligible for a higher loan amount, and therefore the loan ceiling should also be higher. Since $L_{1}$ is always greater than $L_{2}$, we will use $L_{2}$ to determine the loan ceiling, which defines the maximum value of $L_{2}$ to be the loan ceiling.
$L_{2}$ $=\frac{p\underline{Y}}{\frac{2(1+\varepsilon)}{1-(1-e)^{2}}-\delta}$ (2.17)
$=\frac{p\underline{Y}\left[1-(1-e)^{2}\right]}{2(1+\varepsilon)-\delta\left[1-(1-e)^{2}\right]}$
$=\frac{p\underline{Y}\left(2e-e^{2}\right)}{2(1+\varepsilon)-\delta\left(2e-e^{2}\right)}$
Then we take a partial derivative of $L_{2}$ w.r.t. $E$ to obtain
$\frac{\partial L_{2}}{\partial E}$ $=\frac{\partial L_{2}}{\partial e}\cdot\frac{\partial e}{\partial E}$ (2.18)
$=\frac{\left.p\underline{Y}(2-2e)\cdot\left[2(1+\varepsilon)-\delta\left(2e-e^{2}\right)\right]-p\underline{Y}\left(2e-e^{2}\right)(-\delta\right)(2-2e)}{\left[2(1+\varepsilon)-\delta\left(2e-e^{2}\right)\right]^{2}}\times k$
It is clear that the denominator is greater than 0. The numerator is also greater than 0, which can be illustrated by writing out the following result
Numerator $=kp\underline{Y}(2-2e)\times\left[2(1+\varepsilon)-\delta\left(2e-e^{2}\right)+\delta\left(2e-e^{2}\right)\right]$ (2.19)
$=kp\underline{Y}(2-2e)[2(1+\varepsilon)]>0$
A borrower with a higher optimal ESE score, which means he/she has a better overall performance, will be perceived as less risky by the lender. As a result the borrower will be offered a higher loan ceiling, which is the maximum amount of loan that a lender is willing to provide to a borrower. In turn, if a borrower wants to obtain a larger loan amount, he/she may be motivated to improve his/her ESE performance.
2.4 Joint Liability Contracts with a Group Size of $n$
In this subsection we present an individual ESE–joint liability model which formalizes the assumption that the group of borrowers is collectively responsible for the entire group loan (i.e., $nw$). The expected utility of a borrower $i$ who adopts a repayment strategy during any period when they receive financing is specified as
$\max_{E}\text{ }\pi_{i}$ $=\sum_{k=0}^{n-1}\left(\begin{array}[]{c}n-1\\ k\end{array}\right)\cdot\left[e^{n-k}(1-e)^{k}\right]\cdot\left[p\overline{Y}-w-\frac{k(w-p\underline{Y})}{n-k}\right]-C(e)$ (2.20)
$=e(p\overline{Y}-w)-(w-p\underline{Y})\cdot\sum_{k=0}^{n-1}\left(\begin{array}[]{c}n-1\\ k\end{array}\right)\cdot\left[e^{n-k}(1-e)^{k}\right]\cdot\frac{k}{n-k}-C(e)$
$=e(p\overline{Y}-w)-(w-p\underline{Y})\cdot\left((1-e)-(1-e)^{n}\right)-C(e)$
$=ep\overline{Y}-w(1-(1-e)^{n})+p\underline{Y}\left((1-e)-(1-e)^{n}\right)-C(e)$
$=(kE+b)p\overline{Y}-w\left(1-(1-kE-b)^{n}\right)+p\underline{Y}\left[(1-kE-b)-(1-kE-b)^{n}\right]-C(e)$
Constraints similar to those in the case of a group size of two are as follows.
  1. 1.
    There is at least one borrower who is successful at his/her project and capable of paying the entire group’s repayment amount.
    $nw\leqslant p\overline{Y}+(n-1)p\underline{Y}$ (2.21)
  2. 2.
    For each successful borrower, the payoff of strategically defaulting cannot be larger than the payoff of repaying and being refinanced.
    $p\overline{Y}\leqslant p\overline{Y}+(n-1)p\underline{Y}-nw+\delta L$ (2.22)
  3. 3.
    The lender must be able to sustain the lending game, or at least break even, over the periods. Thus the expected repayment amount of each borrower has to be at least as large as $L(1+\varepsilon)$, where $\varepsilon$ is the risk-free rate. It is again important to note that, given that all constraints have been established under the presumption that the group will successfully repay the total loans in order to ensure their eligibility for subsequent loan rounds, we have not factored in the scenario where all borrowers within the group experience failure on the left hand side of this constraint.
    $w\left[1-(1-e)^{n}\right]\geqslant L(1+\varepsilon)$ (2.23)
    Similarly, we obtain $w\geqslant\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$. Thus the value of $w$ should be set as $\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$ due to the negative relationship between $w$ and $\pi_{i}$.
After substituting $w=\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$ in the objective function (which is Equation (2.20)), the ex-ante expected payoff function for the borrower $i$ in a size $n$ joint liability group is represented as
$\max_{E}\pi_{i}=(kE+b)p\overline{Y}-L(1+\varepsilon)+p\underline{Y}\left[(1-kE-b)-(1-kE-b)^{n}\right]-C(e)$ (2.24)
PROPOSITION 4 (Optimal individual–ESE score and group size $n$ have a negative relationship). A higher ESE score indicates a lower default risk, which means that the borrower is more creditworthy and can obtain a loan with a smaller group size. On the other hand each member of the group is jointly responsible for the repayment of the loan, and if one member defaults, the other members will have to repay the outstanding amount. Thus a low-quality borrower may choose to loan with a larger group to reduce the probability of default and minimize the potential burden on any one individual. Next we verify this intuition based on the following derivations. We take a first-order partial derivative of the expected payoff w.r.t $E$ to obtain an optimal individual–ESE score for a borrower
$\frac{\partial\pi}{\partial E}=kp\overline{Y}+p\underline{Y}k\left[n(1-kE-b)^{n-1}-1\right]-k\frac{\partial c(e)}{\partial e}=0$ (2.25)
Since it is hard to find an explicitly analytical form of $n$ and $E$, we adopt an implicit function derivation below. Let $E(n)$ replace $E$ and take the partial derivative w.r.t. $n$ on both sides of $kp\overline{Y}+p\underline{Y}\left[n(1-kE-b)^{n-1}-1\right]-kc(kE+b)=0$. Then we obtain the following result
$p\underline{Y}\left\{(1-kE(n)-b)^{n-1}+n\cdot\frac{\partial\left[e^{(n-1)\log(1-kE(n)-b)}\right]}{\partial n}\right\}-C\left[k\frac{\partial E(n)}{\partial n}\right]=0$ (2.26)
where,
$\frac{\partial\left[e^{(n-1)\log(1-kE(n)-b)}\right]}{\partial n}$ $=e^{(n-1)\log(1-kE(n)-b)}\cdot\left[\frac{(n-1)\cdot(-k)\frac{\partial E(n)}{\partial n}}{1-k\cdot E(n)-b}+\log(1-kE(n)-b)\right]$ (2.27)
$=(1-k(n)-b)^{n-1}\cdot\left[\frac{(n-1)k}{1-kE(n)-b}\cdot\left(-\frac{\partial E(n)}{\partial n}\right)+\log(1-kE(n)-b)\right]$
After reorganizing the formula, we can express the partial derivative of $E(n)$ w.r.t $n$ as
$\frac{\partial E(n)}{\partial n}=\frac{p\underline{Y}\left\{(1-e)^{n-1}[1+n\cdot\log(1-e)]\right\}}{p\underline{Y}\left[n\cdot(1-e)^{n-1}\cdot\frac{(n-1)k}{1-e}\right]+ck}$ (2.28)
It is clear that the denominator in Equation (2.28) is greater than 0. When $n\geqslant 1$ and $e\in(0,1)$, then $n\log(1-e)<-1$, and $(1-e)^{n-1}[1+n\cdot\log(1-e)]<0$. Thus, the numerator is less than 0. As a result $\frac{\partial E(n)}{\partial n}<0$, which proves the optimal individual–ESE score and group size $n$ have a negative relationship. In the light of the negative correlation between optimal individual–ESE scores and a group size, it is recommended that insurance companies prioritize lenders with higher ESE scores when the group size is smaller to mitigate potential risks. Borrowers with high scores are typically perceived to exhibit positive ESE performance, which leads to higher expectations for sustained success and repayment probability over the long term. Despite the relatively small group size, lenders may still have heightened confidence in the ability of high–scoring borrowers to demonstrate honesty and success, and to exhibit willingness to repay their loans. As a result lenders may consider reducing group size requirements for high-scoring borrower groups. Conversely, as the group size increases, the requirement for lenders’ ESE scores may potentially decrease, as a larger group may possess a greater risk mitigation capacity and an increased likelihood of successfully repaying the group loan. Consequently lenders may opt to reduce the score requirements in such cases. PROPOSITION5 (Optimal individual–ESE score converges to a constant level for a large enough group size). Equation (2.28) tells us that $\lim_{n\rightarrow\infty}\frac{\partial E(n)}{\partial n}=0$, which means that the optimal individual–ESE score converges to a constant level when the group size reaches a certain large enough number. To find the limit of the optimal score value, we recall the first-order condition of the ex-ante expected payoff, which is
$\frac{\partial\pi}{\partial E}=kp\overline{Y}+p\underline{Y}k\left[n(1-kE-b)^{n-1}-1\right]-k\frac{\partial C(e)}{\partial e}.$ (2.29)
Similarly we set $C(e)=\frac{1}{2}ce^{2}$, then
$\frac{\partial C(e)}{\partial e}=ce=c(kE+b).$
When $n\rightarrow\infty$, $n(1-kE-b)^{n-1}\rightarrow 0$, then Equation (2.29) approaches $kp(\overline{Y}-\underline{Y})-kc\cdot(kE+b)$. Let $kp(\overline{Y}-\underline{Y})-kc\cdot(kE+b)=0$, the limit of the optimal ESE score will be expressed as a constant
$E\rightarrow\left[\frac{p(\overline{Y}-\underline{Y})}{c}-b\right]/k$ (2.30)
To illustrate this proposition, we provide a numerical example, which is shown in Figure 2. We set the unit price $p=1$, the successful probability $e=k\times E+b=1/100\times E+0$, high yield amount $\overline{Y}=1000$, low yield amount $\underline{Y}=500$, and perform a simulation under the different group size $n\in[1,100]$.
Figure 2: Simulation Results of A Numerical Example
In situations where the group size is small, insurance companies may seek to attract borrowers with high ESE scores in order to mitigate risk. However, as the group size expands, the marginal benefit of an individual effort on group outcomes diminishes, resulting in a reduction of an individual farmer’s optimal effort level. This can create incentives for farmers to contribute less effort, leading to a ‘free-rider’ problem and decreased profits for the group. Therefore the threshold value of the optimal $ESE$ score can serve as a reference point for both the lenders and borrowers, enabling them to avoid an sub-optimal resource allocation and a free-rider problem.
3 Joint Liability Contracts under a Mean–Variance Utility Framework
3.1 The optimal $ESE$ score with different levels of risk aversion
When borrowers consider potential projects, they would like to know how likely it is that they will be able to earn a high return or experience a substantial loss. The variability of profits, or the range of possible returns, is an important consideration for these borrowers. If the potential variability of returns is too high or the risk of substantial losses is too large, the borrower may decide not to invest in the venture. To the best of our knowledge, the current literature on joint liability models focuses only on the expected return or expected profit. This means that they concentrate solely on the central point of an individual’s profit without taking into account the distribution of potential profits or the variability of these returns. However, focusing only on the expected return can be limiting, as it does not account for the risks. In this paper we incorporate the risk (proxied by variance) associated with expected profit into our utility function and conduct further analysis of mean–variance utility in the context of the individual ESE–joint liability model. Specifically the mean–variance utility function for an individual farmer can be incorporated in his/her ex–ante expected profit function in Equation (3.1), where $P$ (a random variable) represents the expected profit and $\gamma$ represents a risk aversion parameter of the farmer
$\max_{E}\pi_{i}$ $=\mathbb{E}(P)-\frac{\gamma}{2}\mathrm{Var}(P)-C(e)$
$=e^{2}(p\overline{Y}-w)+e(1-e)[p\overline{Y}+p\underline{Y}-2w]-C(e)$
$-\frac{\gamma}{2}\left\{e^{2}(p\overline{Y}-w)^{2}+e(1-e)(p\overline{Y}+p\underline{Y}-2w)^{2}\right\}$
$-\frac{\gamma}{2}\left\{-\left[e^{2}(p\overline{Y}-w)+e(1-e)(p\overline{Y}+p\underline{Y}-2w)\right]^{2}\right\}$
$=e[p\overline{Y}+p\underline{Y}-2w]-e^{2}(p\underline{Y}-w)-C(e)$
$-\frac{\gamma}{2}\left\{\left(e^{2}-e^{4}\right)(p\overline{Y}-w)^{2}+\left[e-2e^{2}+2e^{3}-e^{4}\right](p\overline{Y}+p\underline{Y}-2w)^{2}\right\}$
$+\frac{\gamma}{2}\left\{2\left(e^{3}-e^{4}\right)(p\overline{Y}-w)(p\overline{Y}+p\underline{Y}-2w)\right\}$ (3.1)
In order to derive the required optimal individual–ESE score in this setting, we take a first-order partial derivative of the ex-ante profit with respect to the ESE score
$\frac{\partial\pi_{i}}{\partial E}=$ $k(p\overline{Y}+p\underline{Y}-2w)-2ek(p\underline{Y}-w)-k\frac{\partial C(e)}{\partial e}$
-- $\frac{\gamma k}{2}\left\{\left(2e-4e^{3}\right)(p\overline{Y}-w)^{2}+\left(1-4e+6e^{2}-4e^{3}\right)(p\overline{Y}+p\underline{Y}-2w)^{2}\right.$
$\left.-2\left(3e^{2}-4e^{3}\right)(p\overline{Y}-w)(p\overline{Y}+p\underline{Y}-2w)\right\}$ (3.2)
$=$ 00 0
Unfortunately the resulting analytical solution of the ESS score involves a rather complex expression and is difficult to interpret intuitively; so we resort to simulation studies to reveal several insightful patterns. Recall that the probability of achieving a success in a project is given by $e=k\times ESE+b$, wherein the parameters $k$ and $b$ are predetermined. Subsequently the probability may be disaggregated into two constituent components. Specifically the product of $k$ and $ESE$ represents a contribution of a personal impact towards a project success, whereas $b$ represents a constant contribution of exogenous factors beyond the individual control. The value of $k$ is utilized to calibrate the impact of $ESE$ on a project success, while $b$ pertains to the influence of exogenous factors, including climatic and geographical conditions. To determine the value of $b$, an Actuaries Climate Index (ACI) for each region can in principle be utilized as a reference. Briefly the ACI is based on an analysis of six climate indicators, including high and low temperatures, heavy precipitation, drought, and sea level rise. The index measures changes in these indicators over time and across regions, providing indication on how the climate is changing in different parts of a region. It is commonly understood that favorable climate conditions have a positive effect on crop yields, which, in turn, increases the probability of project success. Conversely, unfavorable climate conditions have an adverse impact on the likelihood of success. Hence, we delve deeper into the impact of varying climate conditions caused by climate change on ESE score requirements. By changing the farmer’s degree of risk aversion from 0 to 1 under different cost parameters, a series of simulations are conducted for a high success probability project under favorable climate conditions (where $b=0.7$ in Figure 3), a neutral probability project under neutral climate conditions (where $b=0.5$ in Figure 4), and a relatively low probability project under unfavorable climate conditions (where $b=0.3$ in Figure 5). After observing and comparing these three figures, we conclude our analysis with three findings below. In general any factor that can increase the success probability of the project and the repayment probability can directly or indirectly reduce the requirements of the $ESE$ score. By examining Figures 3, 4, and 5, we notice that, as the probability of success decreases, the $ESE$ score requirement continues to increase if the control cost remains constant; i.e., the probability of a project’s success is inversely proportional to the required ESE score. This is because a higher success rate leads to an increased farmers’ ability to repay their loans, thus lowering the risk of default for the insurance company. As a result the insurance company can lower its $ESE$ score requirements when lending to farmers. From inspecting any one of the three following figures, we see that, as the degree of risk aversion increases, the required ESE score decreases monotonically. This is because an increase in farmers’ risk aversion score leads to a more cautious contract behavior, reducing the probability of a strategic default and increasing the probability of repayment. In other words, a higher degree of risk aversion on the part of the farmers implies a lower level of risk for the insurance company, leading to a lower ESE score requirement. Again, from inspecting any one of the three figures, we can conclude that an increase in the cost parameter will result in a higher requirement for the ESE score. An increase in the cost parameter raises the production expenditure of farmers and burdens their economic resources, thus reducing their profits. Therefore, to avoid a high default probability associated with high costs, the corresponding ESE score requirement should increase. Hence, as the cost parameter increases, the insurance company should raise its ESE score requirement to reduce the risk of default among farmers.
Figure 3: High Successful Probability (Favorable Climate Conditions)
Figure 4: Neutral Successful Probability (Neutral Climate Conditions)
Figure 5: Low Successful Probability (Unfavorable Climate Conditions)
Finally we arrive at the question of how changes in crop yields can impact the required $ESE$ score. To answer this question, we vary the values of yields and observe the effect on the required ESE score. Specifically we consider two scenarios: a ‘Low Yield Case’ where $\overline{Y}=1,000$ and $\underline{Y}=500$, and a ‘High Yield Case’ where $\overline{Y}=600$ and $\underline{Y}=300".$ The results of our simulation in Figure 6 show that there is a negative correlation between crop yields and the required ESE score, i.e., as crop yields increase, the required ESE score decreases. Conversely, as crop yields decrease, the required ESE score increases. When a borrower has lower yields, this can be taken as an indication of a lower income and a potentially higher risk of default; in other words, a higher requirement for borrowers with lower yields is a reflection of the lender’s risk management strategy. By setting stricter requirements for the borrowers with lower yields, the lender is attempting to minimize their exposure to financial risk and ensure that the borrowers are able to repay their loans on time and in full.
Figure 6: High Yields v.s. Low Yields
4 Conclusions
In this paper we have incorporated a novel concept coined as an individual–ESE score into a joint liability model for the purpose of evaluating the sustainability and creditworthiness of individual farmers acting as borrowers in the model. To achieve this objective we initially considered the expected profit of an individual borrower resulting from the specification of his/her utility function. Our primary objective was to determine the optimal group size that maximizes the benefit to borrowers while ensuring that the lender breaks even. Additionally we sought to derive the optimal individual–ESE score, which can provide several benefits to both the lending companies and the farmers as borrowers, such as a better understanding of the borrower’s risk profile, information about the optimal group size for lending, and an improved access to credits. The inclusion of this score in the credit evaluation processes can lead to better–informed lending decisions and sustainable development outcomes. We listed the main results of this paper conveniently in five propositions. Proposition 1 focused on finding an optimal individual-ESE score. Proposition 2 aimed at collapsing two constraints on the loan ceiling to a single constraint. Proposition 3 demonstrated that a higher individual-ESE score leads to a higher loan ceiling. Proposition 4 revealed that the optimal individual-ESE score and group size $n$ have a negative relationship. Finally Proposition 5 showed that the optimal individual–ESE score converges to a constant level when the group size reaches a large enough number. As a next step our study incorporated the risk associated with an expected profit into the utility function of the farmers and analyzed the resulting mean-variance utility framework in the context of the individual $ESE$–joint liability model. The study additionally disentangled the probability of achieving success in a project into two constituent components, namely the impact of personal factors (represented by $k\times ESE$) and exogenous factors beyond individual farmer’s control (represented by $b$), such as climatic and geographical conditions. Our findings reveal that the probability of a project’s success is inversely proportional to the required ESE score. This implies that a higher likelihood of a project success leads to an increased farmers’ ability to repay their loans, thereby lowering the risk of default for the insurance company. As a result the insurance company can lower its $ESE$ score requirements when lending to high-quality farmers. We also found that an increase in farmers’ risk aversion score leads to a more cautious contract behavior, reducing the probability of a strategic default and increasing the probability of repayment. Hence, as the risk–aversion score increases, the corresponding $ESE$ score requirement decreases. Finally an increase in the cost parameter raises the production expenditure of the farmers and places a burden on their economic resources which reduces their profits. Therefore, to avoid a high default probability associated with high costs, the corresponding requirement of an ESE score should be increased. Subsequently we explored the effects of climate change on the required ESE score in order to ensure a successful project implementation. Our simulation results demonstrated that a higher likelihood of a project success, driven by high yields resulting from favorable climate conditions, leads to a lower $ESE$ score requirement in the proposed individual ESE–joint liability model. This suggests that incorporating climate scenarios into the evaluation of creditworthiness can result in better–informed lending decisions and ultimately contribute to more favorable sustainable development outcomes. As an avenue for future research, this study paves the way for an exploration of a partially joint liability’s legal framework which distributes shared responsibility among parties for specific obligations. Unlike a complete joint liability frameworrk, a partially joint liability framework assigns a proportionate debt share to each party. Moreover the concept of a flexible joint liability introduces aspects of innovation. Participants begin with a collaborative lending similar to a standard joint liability contract but, in strategic default scenarios, opt for a punitive phase rather than immediate maximum penalties. After a defined period ($T$ as previously mentioned), defaulters are reintegrated, which offers a path for forgiveness and adherence to the flexible joint liability contract. The diverse joint liability arrangements highlight the need for a thorough and separate analysis. Last but not least, our review of the literature has revealed that while traditional credit metrics concentrate on financial and economic factors, assessing sustainability requires taking into account environmental and social factors as well. Therefore, as a first step of analysis, we have suggested the concept of an individual-ESE score that integrates ESE considerations to offer a more encompassing appraisal of the farming sector. It follows that a potentially most consequential avenue for a future research is to eventually develop a credit scoring system that addresses the unique requirements of the agricultural sector by applying literature–based metrics for agricultural credit scoring systems to the construction of the scores. By compiling and adjusting these metrics, we hope to eventually be able to create a comprehensive credit scoring system with adaptation to climate change for the agricultural sector that addresses limitations of the ESG scores as they are applied to the agricultural sector. This is an important task to take up but is clearly beyond the scope of the present study.